OpenStudy (anonymous):

The triangle shown in subdivided into two additional triangles. All three triangles are similar. Complete and solve the proportion. Simplify radical answers. can some one help???

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

the proportion is 4/h = ?/9

OpenStudy (whpalmer4):

\[\frac{4}{h} = \frac{h}{9}\]

OpenStudy (anonymous):

ik i multiply 4 and 9 but what do i do after that

OpenStudy (whpalmer4):

you can come up with that proportion in at least two different ways: 1) side length 4 to side length h on same triangle 2) side length 4 to side length h on other triangle

OpenStudy (anonymous):

and since they're all 90 degree triangles, you can determine the 3rd side of the 2 smaller triangles and use those to determine the missing 2 sides of the larger triangle by using proportions

OpenStudy (whpalmer4):

cross-multiply: \[4*9 = h*h\]\[36=h^2\]Can you find \(h\)?

OpenStudy (anonymous):

its 6 thankyou

OpenStudy (whpalmer4):

Yes, the value of h is 6. Now can you find the other triangle sides?

OpenStudy (anonymous):

all i needed was h but thakyou though

OpenStudy (whpalmer4):

yes, but can you do it? Maybe you don't need to this time, but there will be a problem in your future where you do :-)

OpenStudy (anonymous):

well maybe i guess lol